# From simplex slicing to sharp reverse Hölder inequalities

### v1

@@ 2.1. Overall strategy and the main idea: $L_{1}$ proxy
$\mathcal{E}_{t}\coloneq X_{1,t}=\mathcal{E}-1-t(\mathcal{E}^{\prime}-1)$
of two-sided exponentials of mean 00. Letting
$\mu_{t}\coloneq\|\mathcal{E}_{t}\|_{1}=\frac{2}{e}\frac{e^{t}}{1+t},$
@@
¯$\bar{\mathcal{E}}_{t}\coloneq\frac{1}{\mu_{t}}\mathcal{E}_{t}.$
@@
Lemma 6 readily reduces the proofs of the main inequalities (6) and (7) to the family {¯$\{\bar{\mathcal{E}}_{t}\}_{t\in[0,1]}$. Indeed, suppose first that $-1<p<0$. Applying (12) twice to $\psi(x)=x^{p}$ on $(0,{+\infty})$ and $\psi(x)=(-x)^{p}$ on $(-\infty,0)$, we get
@@
with the supremum on the left taken over all mean 00 log-concave random variables $X$. By the homogeneity and symmetry of the ratio $\frac{\mathbb{E}|X_{a,b}|^{p}}{(\mathbb{E}|X_{a,b}|)^{p}}$, we can take $10\leq b\leq a=$ 1 which results in
@@
This reduction also follows from Eitan’s results [16] (Theorem 3.1 specialised to $p=1$ therein). His approach uses direct comparison (based on the same main idea of crossing arguments), instead of smoothing. We feel that our Lemma 6 is still of value in this case because of its transparent utilisation of the $L_{1}$ constraint.
Note that ¯$\bar{\mathcal{E}}_{1}=\mathcal{E}_{1}=\mathcal{E}-\mathcal{E}^{\prime}$ has the standard symmetric double-sided exponential density $\frac{1}{2}e^{-|x|}$, consequently |¯$|\bar{\mathcal{E}}_{1}|$ has the same distribution as $\mathcal{E}$, so ¯$\|\bar{\mathcal{E}}_{1}\|_{p}=\Gamma(p+1)^{1/p}$, whereas ¯$\bar{\mathcal{E}}_{0}=\frac{1}{\mu_{0}}(\mathcal{E}-1)=\frac{2}{e}(\mathcal{E}-1)$, hence the right-most equalities in (13) and (14) are as desired.
@@ 2.4. Auxiliary lemmata
Throughout the rest of this paper, we follow the notation introduced in the previous subsection, that is $\mathcal{E},\mathcal{E}^{\prime}$ are i.i.d. $\text{Exp}(1)$, $\mathcal{E}_{t}=\mathcal{E}-1-t(\mathcal{E}^{\prime}-1)$, $\mu_{t}=\mathbb{E}|\mathcal{E}_{t}|$ and ¯$\bar{\mathcal{E}}_{t}=\mu_{t}^{-1}\mathcal{E}_{t}$, for $t\in[0,1]$.
@@
Let $f_{t}\colon[0,{+\infty})\to[0,{+\infty})$ be the density of |¯$|\bar{\mathcal{E}}_{t}|$. For every $10<t<$ 1, each of the functions $f_{t}-f_{0}$ and $f_{1}-f_{t}$ changes sign exactly $3$ times on $(0,{+\infty})$ and has sign pattern (+,,+,)(+,-,+,-)( + , - , + , - ).
@@
When $t=0$ (following the convention in (10) that $e^{-x/t+(1-t)/t}$ is replaced by 00 identically on $(1,{+\infty})$),
@@
The density $f_{t}$ of |¯$|\bar{\mathcal{E}}_{t}|=\mu_{t}^{-1}|\mathcal{E}_{t}|$ is then $f_{t}(x)=\mu_{t}\rho_{t}(\mu_{t}x)$.
Fix $10<t<$ 1. First let us handle the function
@@
Since $\int_{0}^{+\infty}h_{1}=0$, and $h_{1}$ is not identically 00, it must change sign at least once; additionally $\int_{0}^{+\infty}h_{1}(x)x\mathrm{d}x=0$, so $h_{1}$ need change sign at least twice on $(0,{+\infty})$, by Lemma 5. It is evident that the ratio $r_{1}(x)=f_{t}(x)/f_{1}(x)=\mu_{t}\rho_{t}(\mu_{t}x)e^{x}$ is convex on each of the intervals $[0,\alpha_{t}]$ and $[\alpha_{t},{+\infty})$, where $\alpha_{t}\coloneq\mu_{t}^{-1}(1-t)$ denotes the point where the formula defining $f_{t}(x)$ changes. Moreover, $r_{1}(x)$ is strictly increasing on $[0,\alpha_{t}]$ (as the product of two strictly increasing functions) with $r_{1}(0)=\mu_{t}\rho_{t}(0)=\mu_{t}^{2}<1$. It is also clear that $r_{1}(x)\to{+\infty}$ as $x\to{+\infty}$. If $r_{1}(\alpha_{t})\leq 1$, then $r_{1}(x)=1$ would have no solutions on $(0,\alpha_{t})$ and exactly $1$ solution on $(\alpha_{t},{+\infty})$, resulting in $f_{1}-f_{t}$ having at most one sign change, ruled out earlier. Therefore, $r_{1}(\alpha_{t})>1$. Consequently, $r_{1}(x)=1$ has exactly one solution on $(0,\alpha_{t})$ (by monotonicity) and exactly $2$ solutions on $(\alpha_{t},{+\infty})$ (by convexity), resulting in $f_{1}-f_{t}$ having exactly three sign changes and the sign pattern (+,,+,)(+,-,+,-)( + , - , + , - ) (because $r_{1}(0)<1$).
@@
We will frequently use that $\mu_{t}>\mu_{0}$. Recall $\rho_{0}(\mu_{0}x)$ has a jump discontinuity at $x_{0}\coloneq\mu_{0}^{-1}$ and the formula for $\rho_{t}(\mu_{t}x)$ changes at $x=\alpha_{t}=\mu_{t}^{-1}(1-t)<\mu_{0}^{-1}=x_{0}$. It will be convenient to split the analysis into the consecutive intervals: $[0,\alpha_{t}]$, $(\alpha_{t},x_{0}]$ and $(x_{0},+\infty)$.
On $[0,\alpha_{t}]$, we have $\frac{1}{2}h_{0}(x)=\mu_{t}\frac{e^{t-1}}{1+t}\cosh(\mu_{t}x)-\mu_{0}e^{-1}\cosh(\mu_{0}x)$. Since $\mu_{t}>\mu_{0}$ and $\frac{e^{t-1}}{1+t}>e^{-1}$, we have $h_{0}>0$ on $[0,\alpha_{t}]$.
On $[\alpha_{t},x_{0}]$, $f_{t}$ is now strictly decreasing, whereas $f_{0}(x)=\mu_{0}^{2}\cosh(\mu_{0}x)$ is still strictly increasing, so $h_{0}$ changes sign at most once on this interval.
@@
Function $h_{0}$ admits one more sign change, at the discontinuity point $x_{0}$, so at most $3$ sign changes on $(0,{+\infty})$, and minimum $2$ sign changes. As remarked, near 00, $h_{0}$ is positive, and eventually negative, thus it need change sign an odd number of times. Thus we get exactly $3$ sign changes. Since $f_{t}(x)<f_{0}(x)$ as $x\to{+\infty}$, the sign pattern is (+,,+,)(+,-,+,-)( + , - , + , - ).
@@
Recall $\mathcal{E}_{t}=\mathcal{E}-t\mathcal{E}^{\prime}-(1-t)$. Using that the distribution of $\mathcal{E}-t\mathcal{E}^{\prime}$ is the same as a mixture of two exponentials: $\mathcal{E}$ with probability $\frac{1}{1+t}$, and $-t\mathcal{E}$ with probability $\frac{t}{1+t}$, we have
@@
    The function $t\mapsto\|\bar{\mathcal{E}}_{t}\|_{3}$ is unimodal on $[0,1]$: first strictly decreasing, then strictly increasing, with ¯$\|\bar{\mathcal{E}}_{t}\|_{3}\leq\|\bar{\mathcal{E}}_{1}\|_{3}$ for all $\frac{1}{5}\leq t\leq 1$.
@@
¯$\|\bar{\mathcal{E}}_{t}\|_{2}^{2}=\mu_{t}^{-2}\operatorname{Var}(\mathcal{E}_{t})=\mu_{t}^{-2}(1+t^{2})=\frac{1}{4}e^{2(1-t)}(1+t^{2})(1+t)^{2}.$
@@
Its sign is the same as the sign of the function in the square parentheses $h(t)=2t^{2}+2t+1-t^{4}-4e^{t-1}$ which we now analyse. Since $h^{\prime\prime\prime}(t)=-24t-4e^{t-1}<0$, $h^{\prime}$ is concave. Moreover, $h^{\prime}(0)=2-4/e>0$ and $h^{\prime}(1)=-2<0$, so $h^{\prime}$ is first positive then negative on $[0,1]$. Consequently, $h$ first increases then decreases. Since $h(0)=1-4/e<0$ and $h(1)=0$, we conclude that $h$ is first negative, then positive on $(0,1)$ with exactly one zero. As a result $t\mapsto\mathbb{E}|\bar{\mathcal{E}}_{t}|^{3}$ is first decreasing then increasing. Finally, we check that $\mathbb{E}|\bar{\mathcal{E}}_{1/5}|^{3}=5.97..<6=\mathbb{E}|\bar{\mathcal{E}}_{1}|^{3}$ which finishes the proof.
@@
As in the proof of Lemma 6, we will repeatedly argue that certain integral inequalities hold by rearranging the integrand to show that it is pointwise nonnegative, analysing sign patterns of sums of certain power functions. We record the following elementary lemma, adapted to concrete functions stemming from our setting.
@@
and hence $g(x^{\prime})$ is nonzero and has the same sign as $(0-p)(1-p)(q-p)$. This holds for any $x^{\prime}\in(0,x_{1})$. Similar reasoning in the cases when $x^{\prime}\in(x_{1},x_{2})$ or $(x_{2},x_{3})$ or $(x_{3},+\infty)$ yields the desired result.∎
@@ 2.5. Proof of (13)
Fix $10<t<$ 1. First note that thanks to Lemma 9 (i) and (iii), the function
@@
is positive at $q=2$ and negative at $q=4$, so by its continuity, there is $q=q(t)\in(2,4)$ where it vanishes. We fix such a value and call it $q$. As in Lemma 7, let $f_{t}$ be the density of |¯$|\bar{\mathcal{E}}_{t}|$. Fix nonzero $p\in(-1,1)$. Observe that for arbitrary $\alpha,\beta,\gamma\in\mathbb{R}$, we can write
@@
as each of the integrals against $\alpha$, $\beta x$ and $\gamma x^{q}$ is zero. We know from Lemma 7 that the first parenthesis has exactly $3$ sign changes on $(0,{+\infty})$, say at $x_{1}<x_{2}<x_{3}$, with sign pattern (+,,+,)(+,-,+,-)( + , - , + , - ). We choose $\alpha,\beta,\gamma$ from Lemma 10 for these nodes. Moreover, from the lemma, the sign pattern of the second parenthesis is (+,,+,)(+,-,+,-)( + , - , + , - ) when $-1<p<0$, in which case the integrand is pointwise nonnegative and we get
@@
that is ¯$\|\bar{\mathcal{E}}_{t}\|_{p}\geq\|\bar{\mathcal{E}}_{1}\|_{p}$ ($p$ is negative). Similarly, when $10<p<$ 1, the sign pattern is flipped resulting with the reverse inequality above and consequently again ¯$\|\bar{\mathcal{E}}_{t}\|_{p}\geq\|\bar{\mathcal{E}}_{1}\|_{p}$, as desired. ∎

### v2

@@ 2.1. Overall strategy and the main idea: $L_{1}$ proxy
$\mathcal{E}_{t}=X_{1,t}=\mathcal{E}-1-t(\mathcal{E}^{\prime}-1)$
of two-sided exponentials of mean 0. Letting
$\mu_{t}=\|\mathcal{E}_{t}\|_{1}=\frac{2}{e}\frac{e^{t}}{1+t},$
@@
$\bar{\mathcal{E}}_{t}=\frac{1}{\mu_{t}}\mathcal{E}_{t}.$
@@
Lemma 6 readily reduces the proofs of the main inequalities (6) and (7) to the family $\{\bar{\mathcal{E}}_{t}\}_{t\in[0,1]}$. Indeed, suppose first that 1<p<0-1<p<0. Applying (12) twice to $\psi(x)=x^{p}$ on $(0,{+\infty})$ and $\psi(x)=(-x)^{p}$ on $(-\infty,0)$, we get
@@
with the supremum on the left taken over all mean 0 log-concave random variables $X$. By the homogeneity and symmetry of the ratio $\frac{\mathbb{E}|X_{a,b}|^{p}}{(\mathbb{E}|X_{a,b}|)^{p}}$, we can take $0\leq b\leq a=1$ which results in
@@
Conceptually, one might therefore think of Lemma 6 as a localisation step, in that it reduces the optimisation problem at hand over all (constrained) log-concave densities to simple concrete densities (two-sided exponentials). As discussed earlier, the off-hand localisation method of Fradelizi and Guedon from [20] does not seem to be a feasible approach as the resulting problem involves too many parameters to handle. Finally, as another point of comparison, Eitan’s main argument from [17] leads to the same reduction (Theorem 3.1 specialised to $p=1$ therein). His approach also relies on crossing arguments, but via direct comparison, viz. bounding the number of zeros of polynomial-type functions (Chebyshev’s systems). Our application of Lemma 6 on the other hand strongly relies on the convexity of $|x|^{p}$ as well as the interplay between the mean zero assumption and the $L_{1}$ constraint, which we have strived to make explicit by phrasing Lemma 6 in a more robust setting concerning the optimisation of integrals against convex functionals $\psi$.
It is also worth comparing Lemma 6 to Fradelizi’s results from [19], in particular Theorem 8, and the upper bound of (6), therein, which states that among all log-concave probability densities $f$ of mean 0, with a fixed value at 0, the symmetric exponential density maximises the integral against any even convex function. We can recover Fradelizi’s upper bound using Lemma 6 (nota bene, the assumption that $\phi$ is even is not needed). We state this as the following corollary to Lemma 6, and sketch the argument.
Corollary 7 (Upper bound (6) in [19]).
Let $\phi\colon\mathbb{R}\to\mathbb{R}$ be a convex function. For every log-concave probability density function $f\colon\mathbb{R}\to[0,+\infty)$ with mean 0, that is $\int_{-\infty}^{\infty}xf(x)\mathrm{d}x=0$, we have
$\int_{-\infty}^{\infty}\phi(x)f(x)\mathrm{d}x\leq\int_{-\infty}^{\infty}\phi(x)\cdot\Big(f(0)e^{-2f(0)|x|}\Big)\mathrm{d}x.$
Sketch.
We apply (12) twice, to $\psi(x)=\phi(\lambda x)$, $\lambda=\int|x|f(x)$, restricted to $(0,+\infty)$ and $(-\infty,0)$, and obtain $\int\phi\cdot g\geq\int\phi\cdot f$, where $g=g_{a,b}$ is the two-sided exponential density (10), matching the quantities $\int_{0}^{\infty}f$ and $\int_{0}^{\infty}tf(t)\mathrm{d}t$, (instead of $f(0)$). We then consider $\tilde{g}(x)=g(0)\exp\{-2g(0)|x|\}$ and argue that $\tilde{g}-g$ has sign pattern (+,,+)(+,-,+). Since $\phi$ is convex, and both densities $g$ and $\tilde{g}$ have mean 0, this gives $\int\phi\cdot\tilde{g}\geq\int\phi\cdot g$. Now we consider $\tilde{f}(x)=f(0)\exp\{-2f(0)|x|\}$. Crucially, it can be showed that $g(0)\geq f(0)$ (assuming without loss of generality, $a\geq b>0$, the calculations in the proof of Lemma 6 give $g(0)=\frac{1}{a}\frac{1}{e}\frac{e^{b/a}}{1+b/a}=\frac{(\int_{0}^{\infty}f)^{2}}{\int_{0}^{\infty}xf(x)}$ which is at least $f(0)$, e.g. by Lemma 2 in [41]). Thus, equivalently, $\tilde{f}-\tilde{g}$ has sign patter (+,,+)(+,-,+), and, as a result, $\int\phi\cdot\tilde{f}\geq\int\phi\cdot\tilde{g}$, leading to $\int\phi\cdot\tilde{f}\geq\int\phi\cdot f$, as desired. ∎
Note that $\bar{\mathcal{E}}_{1}=\mathcal{E}_{1}=\mathcal{E}$-$\mathcal{E}^{\prime}$ has the standard symmetric double-sided exponential density $\frac{1}{2}e^{-|x|}$, consequently $|\bar{\mathcal{E}}_{1}|$ has the same distribution as $\mathcal{E}$, so $\|\bar{\mathcal{E}}_{1}\|_{p}=\Gamma(p+1)^{1/p}$, whereas $\bar{\mathcal{E}}_{0}=\frac{1}{\mu_{0}}(\mathcal{E}-1)=\frac{2}{e}(\mathcal{E}-1)$, hence the right-most equalities in (13) and (14) are as desired.
@@ 2.4. Auxiliary lemmata
Throughout the rest of this paper, we follow the notation introduced in the previous subsection, that is $\mathcal{E},\mathcal{E}^{\prime}$ are i.i.d. $\text{Exp}(1)$, $\mathcal{E}_{t}=\mathcal{E}-1-t(\mathcal{E}^{\prime}-1)$, $\mu_{t}=\mathbb{E}|\mathcal{E}_{t}|$ and $\bar{\mathcal{E}}_{t}=\mu_{t}^{-1}\mathcal{E}_{t}$, for $t\in[0,1]$.
@@
Let $f_{t}\colon[0,{+\infty})\to[0,{+\infty})$ be the density of $|\bar{\mathcal{E}}_{t}|$. For every $0<t<1$, each of the functions $f_{t}-f_{0}$ and $f_{1}-f_{t}$ changes sign exactly $3$ times on $(0,{+\infty})$ and has sign pattern (+,,+,)(+,-,+,-).
@@
When $t=0$ (following the convention in (10) that $e^{-x/t+(1-t)/t}$ is replaced by 0 identically on $(1,{+\infty})$),
@@
The density $f_{t}$ of $|\bar{\mathcal{E}}_{t}|=\mu_{t}^{-1}|\mathcal{E}_{t}|$ is then $f_{t}(x)=\mu_{t}\rho_{t}(\mu_{t}x)$.
Fix $0<t<1$. First let us handle the function
@@
Since $\int_{0}^{+\infty}h_{1}=0$, and $h_{1}$ is not identically 0, it must change sign at least once; additionally $\int_{0}^{+\infty}h_{1}(x)x\mathrm{d}x=0$, so $h_{1}$ need change sign at least twice on $(0,{+\infty})$, by Lemma 5. It is evident that the ratio $r_{1}(x)=f_{t}(x)/f_{1}(x)=\mu_{t}\rho_{t}(\mu_{t}x)e^{x}$ is convex on each of the intervals $0,\alpha_{t}]$ and $\alpha_{t},{+\infty})$, where $\alpha_{t}=\mu_{t}^{-1}(1-t)$ denotes the point where the formula defining $f_{t}(x)$ changes. Moreover, $r_{1}(x)$ is strictly increasing on $0,\alpha_{t}]$ (as the product of two strictly increasing functions) with $r_{1}(0)=\mu_{t}\rho_{t}(0)=\mu_{t}^{2}<1$. It is also clear that $r_{1}(x)\to{+\infty}$ as $x\to{+\infty}$. If $r_{1}(\alpha_{t})\leq 1$, then $r_{1}(x)=1$ would have no solutions on $(0,\alpha_{t})$ and exactly $1$ solution on $(\alpha_{t},{+\infty})$, resulting in $f_{1}-f_{t}$ having at most one sign change, ruled out earlier. Therefore, $r_{1}(\alpha_{t})>1$. Consequently, $r_{1}(x)=1$ has exactly one solution on $(0,\alpha_{t})$ (by monotonicity) and exactly $2$ solutions on $(\alpha_{t},{+\infty})$ (by convexity), resulting in $f_{1}-f_{t}$ having exactly three sign changes and the sign pattern (+,,+,)(+,-,+,-) (because $r_{1}(0)<1$).
@@
We will frequently use that $\mu_{t}>\mu_{0}$. Recall $\rho_{0}(\mu_{0}x)$ has a jump discontinuity at $x_{0}=\mu_{0}^{-1}$ and the formula for $\rho_{t}(\mu_{t}x)$ changes at $x=\alpha_{t}=\mu_{t}^{-1}(1-t)<\mu_{0}^{-1}=x_{0}$. It will be convenient to split the analysis into the consecutive intervals: $0,\alpha_{t}]$, $(\alpha_{t},x_{0}]$ and $(x_{0},+\infty)$.
On $0,\alpha_{t}]$, we have $\frac{1}{2}h_{0}(x)=\mu_{t}\frac{e^{t-1}}{1+t}\cosh(\mu_{t}x)-\mu_{0}e^{-1}\cosh(\mu_{0}x)$. Since $\mu_{t}>\mu_{0}$ and $\frac{e^{t-1}}{1+t}>e^{-1}$, we have $h_{0}>0$ on $0,\alpha_{t}]$.
On $\alpha_{t},x_{0}]$, $f_{t}$ is now strictly decreasing, whereas $f_{0}(x)=\mu_{0}^{2}\cosh(\mu_{0}x)$ is still strictly increasing, so $h_{0}$ changes sign at most once on this interval.
@@
Function $h_{0}$ admits one more sign change, at the discontinuity point $x_{0}$, so at most $3$ sign changes on $(0,{+\infty})$, and minimum $2$ sign changes. As remarked, near 0, $h_{0}$ is positive, and eventually negative, thus it needs to change sign an odd number of times. Thus we get exactly $3$ sign changes. Since $f_{t}(x)<f_{0}(x)$ as $x\to{+\infty}$, the sign pattern is (+,,+,)(+,-,+,-).
@@
Recall $\mathcal{E}_{t}=\mathcal{E}-t\mathcal{E}^{\prime}-(1-t)$. Using that the distribution of $\mathcal{E}-t\mathcal{E}^{\prime}$ is the same as a mixture of two exponentials: $\mathcal{E}$ with probability $\frac{1}{1+t}$, and $t\mathcal{E}$ with probability $\frac{t}{1+t}$, we have
@@
    The function $t\mapsto\|\bar{\mathcal{E}}_{t}\|_{3}$ is unimodal on $[0,1]$: first strictly decreasing, then strictly increasing, with $\|\bar{\mathcal{E}}_{t}\|_{3}\leq\|\bar{\mathcal{E}}_{1}\|_{3}$ for all $\frac{1}{5}\leq t\leq 1$.
@@
$\|\bar{\mathcal{E}}_{t}\|_{2}^{2}=\mu_{t}^{-2}\operatorname{Var}(\mathcal{E}_{t})=\mu_{t}^{-2}(1+t^{2})=\frac{1}{4}e^{2(1-t)}(1+t^{2})(1+t)^{2}.$
@@
Its sign is the same as the sign of the function in the square parentheses $h(t)=2t^{2}+2t+1-t^{4}-4e^{t-1}$ which we now analyse. Since $^{\prime\prime\prime}(t)=-24t-4e^{t-1}<0, hh^{\prime}$ is concave. Moreover, $h^{\prime}(0)=2-4/e>0$ and $h^{\prime}(1)=-2<0$, so $h^{\prime}$ is first positive then negative on $[0,1]$. Consequently, $h$ first increases then decreases. Since h(0)=14/e<0h(0)=1-4/e<0 and $h(1)=0$, we conclude that $h$ is first negative, then positive on $(0,1)$ with exactly one zero. As a result $t\mapsto\mathbb{E}|\bar{\mathcal{E}}_{t}|^{3}$ is first decreasing then increasing. Finally, we check that $\mathbb{E}|\bar{\mathcal{E}}_{1/5}|^{3}=5.97..<6=\mathbb{E}|\bar{\mathcal{E}}_{1}|^{3}$ which finishes the proof.
@@
As in the proof of Lemma 6, we will repeatedly argue that certain integral inequalities hold by rearranging the integrand to show that it is pointwise nonnegative, analysing sign patterns of sums of certain power functions. We record the following elementary lemma, adapted to concrete functions stemming from our setting. Using an exponential change of variables, the following can be found as the $n=3$ case of [18, Lemma A.2].
@@
and hence $g(x^{\prime})$ is nonzero and has the same sign as (0p)(1p)(qp)(0-p)(1-p)(q-p). This holds for any $x^{\prime}\in(0,x_{1})$. Similar reasoning in the cases when $x^{\prime}\in(x_{1},x_{2})$ or $(x_{2},x_{3})$ or $(x_{3},+\infty)$ yields the desired result.∎
@@ 2.5. Proof of (13)
Fix $0<t<1$. First note that thanks to Lemma 10 (i) and (iii), the function
@@
is positive at $q=2$ and negative at $q=4$, so by its continuity, there is $q=q(t)\in(2,4)$ where it vanishes. We fix such a value and call it $q$. As in Lemma 8, let $f_{t}$ be the density of $|\bar{\mathcal{E}}_{t}|$. Fix nonzero $p\in(-1,1)$. Observe that for arbitrary $\alpha,\beta,\gamma\in\mathbb{R}$, we can write
@@
as each of the integrals against $\alpha$, $\beta x$ and $\gamma x^{q}$ is zero. We know from Lemma 8 that the first parenthesis has exactly $3$ sign changes on $(0,{+\infty})$, say at $x_{1}<x_{2}<x_{3}$, with sign pattern (+,,+,)(+,-,+,-). We choose $\alpha,\beta,\gamma$ from Lemma 11 for these nodes. Moreover, from the lemma, the sign pattern of the second parenthesis is (+,,+,)(+,-,+,-) when 1<p<0-1<p<0, in which case the integrand is pointwise nonnegative and we get
@@
that is $\|\bar{\mathcal{E}}_{t}\|_{p}\geq\|\bar{\mathcal{E}}_{1}\|_{p}$ ($p$ is negative). Similarly, when $0<p<1$, the sign pattern is flipped resulting with the reverse inequality above and consequently again $\|\bar{\mathcal{E}}_{t}\|_{p}\geq\|\bar{\mathcal{E}}_{1}\|_{p}$, as desired. ∎